OpenStudy (anonymous):

simplify by factoring 4√243x^12y^14

OpenStudy (anonymous):

Have you tried anything?

OpenStudy (anonymous):

yes but i am lost with these square root I can follow a pattern but there is not a pattern for this one in my manual

OpenStudy (anonymous):

no x^12 and y^14

OpenStudy (mertsj):

\[4\sqrt{243x ^{12}y ^{14}}\]

OpenStudy (mertsj):

Do I have the problem right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[ 36 \sqrt{3} \sqrt{x^{12} y^{14}} \]

OpenStudy (mertsj):

\[4\sqrt{3*3*3*3*x ^{6}*x ^{6}*y ^{7}*y ^{7}}\]

OpenStudy (mertsj):

\[4(3)(3)(x ^{6}y ^{7})= 36x ^{6}y ^{7}\]

OpenStudy (turingtest):

you dropped a 3...

OpenStudy (mertsj):

Yep you are correct so....

OpenStudy (anonymous):

and finallly,\[ \huge 36 \sqrt{3} \sqrt{x^{12} y^{14}} = 36 x^6 y^7 \sqrt{3} \]

OpenStudy (mertsj):

\[36\sqrt{3} x ^{6}y ^{7}\]

OpenStudy (anonymous):

which one is correct

OpenStudy (mertsj):

They are both the same. The sqrt3 is just moved

OpenStudy (anonymous):

Of-course the one I posted ...

OpenStudy (turingtest):

FFM isn't being sincere they are both right above

OpenStudy (anonymous):

I am always sincere.

OpenStudy (mertsj):

Except when you aren't.

OpenStudy (turingtest):

why must you confuse the student so? Mertsj is right.

OpenStudy (turingtest):

...that they are both right lol

OpenStudy (anonymous):

I am not confusing anyone, I never said anything about the other answer.

OpenStudy (turingtest):

fair enough

OpenStudy (mertsj):

Oh stop it. You guys clearly don't have enough to do.

OpenStudy (anonymous):

i thank you for your help i clearly need to study this much more i envy all of you

OpenStudy (mertsj):

You are welcome

OpenStudy (anonymous):

Glad to help :)